Please assist with how to figure these out with excel??

You’ve decided you want to get your MBA, which will cost you $15,000. If you have $10,000 now and can make 7.5% on your money, how soon can you attend classes?

You’ve just been hired on your first job! Congratulations. The company has a 401(k) plan and you’ve decided to put 10% of your $45,000 salary away each year. If you can earn 8% annually on your money, how much will you have 35 years from now when you’re ready to retire?

You’re aunt Edna just retired at the age of 65 years old. She has an annuity with $750,000 and wants to begin taking payments until she’s 90 years old. If the annuity can earn 4.5% annually, how much can she take each year before the account gets to $0?

You have to choose between investing your $1000 in a band CD that pays 2.5% annually, or into a 20 year treasury bond that pays 2.2% semi-annually. Which investment pays a better rate of return?
Answer #1

1.
=NPER(7.5%,0,-10000,15000)=5.60649058307827

2.
=FV(8%,35,-45000*10%)=775425.616555532

3.
=PMT(4.5%,25,-750000)=50579.2710290209

4.
EAR of CD=2.5%
EAR of semiannual bond=(1+2.2%/2)^2-1=2.2121%

CD annual pays better return